### v2

# Ad-hoc Concept Forming in the Game Codenames as a Means for Evaluating Large Language Models

Abstract
This study utilizes the game Codenames as a benchmarking tool to evaluate large language models (LLMs) with respect to specific linguistic and cognitive skills. LLMs play each side of the game, where one side generates a clue word covering several target words and the other guesses those target words. We designed various experiments by controlling the choice of words (abstract vs. concrete words, ambiguous vs. monosemic) or the opponent (programmed to be faster or slower in revealing words). Recent commercial and open-weight models were compared side-by-side to find out factors affecting their performance. The evaluation reveals details about their strategies, challenging cases, and limitations of LLMs.
Ad-hoc Concept Forming in the Game Codenames
as a Means for Evaluating Large Language Models
Sherzod Hakimov1, Lara Pfennigschmidt1, David Schlangen1,2 1Computational Linguistics, Department of Linguistics University of Potsdam, Germany 2German Research Center for Artificial Intelligence (DFKI), Berlin, Germany firstname.lastname@uni-potsdam.de
1 Introduction
The astounding abilities of large language models (LLMs) have led to what could be called a ‘crisis of evaluation’, where the previous paradigm of evaluating natural language processing (NLP) models—through pairs of problem instance and expected response—does not fit well any more. First, the main mode of usage of LLMs is through their embedding in a “chatbot”, often across multiple turns, which is not represented by the reference-based evaluation mode. Second, the closed nature and sheer size of the training data, often acquired through automatic means from the open internet, raises fears that the usual test datasets have been ingested and hence the training data has become contaminated, rendering the value of the tests even more doubtful Golchin and Surdeanu (2024); Deng et al. (2024).
Figure 1: Overview of the proposed approach where an LLM poses as both Spymaster (clue giver) and Field Operative (guesser) and plays against a mock opponent. The GameMaster orchestrates the game play by keeping and updating the game state on the board (image from https://codenames.game/.
The use of games as an interactive environment where LLMs are tasked to perform certain actions and scored whether they achieve the task or not has emerged as a response to this situation Chalamalasetti et al. (2023); Wu et al. (2024); Zhou et al. (2024), allowing the evaluation to more closely approximate the interactive use situation, and overcoming the data contamination problem in two ways. First, even for already documented games, it is easy to create new instances that lead to game play that differs from what is in existing data Beyer et al. (2024). Second, to further extend the range of evaluated phenomena, it is only necessary to add new game implementations, rather than to create new datasets. It is this second dimension that we explore in this paper.
We have implemented the game Codenames  [^1] Source code (”codenames” directory): https://github.com/clembench/clembench as a challenging means (already investigated in psycholinguistic studies, see below) for evaluating certain language-use capabilities. In this game, a first player needs to provide to a second player a clue which singles out certain words within a larger set of words given to both players (see Figure 6 below for examples).
The game requires cooperation, with the first player needing to form and name an ad-hoc concept that spans the target concepts, in a way that they assume is understandable to others (theory of mind) (Kim et al., 2019). Players need to connect words in a wide variety of relations such as homonyms, antonyms, rhymes, or popular culture references (Jaramillo et al., 2020). Clue generation is also a task of co-creativity (Spendlove and Ventura, 2023), testing skills in the evaluation of semantic relatedness of words and common-sense reasoning (Bitton et al., 2022) as well as the ability to constrain clues and negatively associate them with any non-team word.
Players need to predict the partners’ behaviour and knowledge (Cserháti et al., 2022; Kumar et al., 2021), so one cannot simply optimise their own behaviour (Jaramillo et al., 2020) without acknowledging the cultural background and knowledge level of their teammates (Shaikh et al., 2023), hence requiring cooperation. Figure 1 shows the overall idea where players (Spymaster and Field Operative) are LLMs that play against a programmed mock opponent. The programmatic GameMaster component (part of the framework we use) orchestrates the game play by providing inputs and generated outputs among the parties, checks whether players comply to the rules of the game.
Our contributions are as follows: i) benchmarking LLMs to test their ad-hoc concept generation, cooperation, pragmatic reasoning capabilities, ii) comparison of open-weight and commercial models under various experiments, iii) in-depth analysis of how best-performing models navigate the task.
2 Related Work
Earlier work focused on using various word embedding techniques (choose the clue that is closest to targets and most distant to distractors) Kim et al. (2019); Jaramillo et al. (2020). Later, such methods were combined with associative methods that use language graphs for generating clues or guessing Koyyalagunta et al. (2021). Other approaches involve concentrating on word co-occurrence measurements  de Rijk and Marecek (2020); Cserháti et al. (2022) for capturing synonymy, semantic similarity, or word-relatedness measure instead of just focusing on word embeddings.
Later research started looking into using LLMs to generate clues Spendlove and Ventura (2022). The idea of benchmarking LLMs led to the development of various datasets, e.g. BigBench Srivastava et al. (2022) includes Codenames as one of the many tasks to test emergent abilities of models Wei et al. (2022); Ozturkler et al. (2023); Lu et al. (2023). Stephenson et al. (2024) recently explored using Codenames to benchmark LLMs where two pairs of LLMs (red vs. blue team) play against each other. Our method differs from theirs by comparing the language model team against a deterministic opponent. Not using a deterministic opponent could lead to different results every time the same game is played due to the non-deterministic nature of language models Song et al. (2024). Another extension of our work lies in the experimental setup, where we study the effect of selecting words on a board and their relations in much more depth.
3 The game: Codenames
Codenames (Chvátil and Kučerovskỳ, 2015) is a cooperative board game with two teams (blue and red) that try to uncover their team agents’ code names before the other team finds all of theirs. The board is set up with 25 word cards. Each team has a “Spymaster” that knows which words on the board represent their team (8+1 for the starting team), the opposing team (8), innocent bystanders (7), and the assassin (1). The team starting the game has one more word to uncover to balance out the advantage of going first. Our implemented version deploys one Spymaster and Field Operative on the same team. The opponent team is mocked with an ideal behaviour of revealing n own words each turn.
The Spymaster takes turns providing clues for their teammates – the “Field Operatives”. Each clue consists of a word related to one or more (code names) targeted words. It has to output in this way:
    CLUE: <clue>
    TARGETS: <list of targets>
Only the clue is passed on to the Field Operative, who then guesses the matching words:
    GUESS: <list of guesses>
If the guess is correct, the team can continue guessing as many names as the Spymaster indicated in their clue. If the team is unsure, they can also end their turn voluntarily. If the team’s guess is incorrect, meaning they contacted an innocent bystander or an word of the opposing team, the identity is revealed, and the team’s turn ends. If the team uncovers any assassin word, the team immediately loses the game.
We have implemented the game using the clembench (Chalamalasetti et al., 2023) framework where the GameMaster orchestrates the gameplay by 1) checking the required formatting of generated outputs by Spymaster or Field Operative, 2) passing the outputs between players. The Spymaster and Field Operative prompts are given in Appendix 9.
4 Experimental Setup
4.1 Board Generation
We used different sets of words to design experiments. Each experiment includes 10 instances (boards) where the words are chosen randomly from a specific set. The default mock opponent uncovers one word per turn (n=1). The default word list is by de Rijk (2020) with one assassin word per board. We defined the following experiments by changing specific default parameters, which correspond to 130 instances:
$\bullet$ Risk level: We included five assassin words in the set called high risk. The low risk set has no assassin words. The rationale here is to see whether models target less number of words to mitigate the risk of revealing assassin words.
$\bullet$ Word association: We selected 45 category norms (e.g. bird name, kitchen utensil, country, military title, etc.) from the corpus by Castro et al. (2021). The easy set is created by selecting 3-5 categories, sampling words for each category, and assigning them to the same team (3-5 turns by targeting the category). The difficult set is created by ensuring that sampled words are distributed across all possible groups (team, opponent, innocent, assassin) and not assigned to the same team. The rationale here is whether models actually can capture those obvious associations on the easy set and whether they can play the difficult one at all.
$\bullet$ Opponent level: We created three sets where the mock opponent turns two, one or none words per turn, which correspond to difficult, easy, and none levels, respectively. The rationale here is to check whether LLMs can play against a faster opponent that constantly reveals two words at a time.
$\bullet$ Word frequency: All nouns from the SUBTLEX-US corpus Brysbaert and New (2009) were filtered out to create two sets for low and high-frequency lists. We used the top and bottom 250 words for the frequency lists of the high and low. Typical human players would usually struggle with low frequency words and our rationale is to check whether it poses a similar challenge to LLMs too.
$\bullet$ Word ambiguity: The corpus provided by Beekhuizen et al. (2021) includes monosemes (words with single sense) and homonyms (words with multiple senses). The ambiguous set is composed of homonyms while the unambiguous one includes the monosemes. The hypothesis here is that words with multiple meanings are easier to find connections between them than ambiguous words.
$\bullet$ Word concreteness: Two sets of words where one corresponds to concrete concepts and the other includes abstract ones. Brysbaert et al. (2014) collected word concreteness ratings (Likert scale between 1-5). We used the top 500 words with the lowest and highest concreteness ratings for abstract and concrete word lists, respectively. The hypothesis here to check whether LLMs play better with concrete words as it is easier for human players to find association between them in contrast to abstract concepts.
4.2 Metrics
The clembench framewor measures how many of the instances (boards) have resulted in a Played or Aborted state. The gameplay is marked as Aborted if either player does not follow the formatting instructions when generating an output (as explained in Section 3). Played is the ratio of remaining gameplays (episodes) where formatting instructions have been followed. The Played ratio is further divided into Success if the team reveals own words faster than mock opponent, or Lose if an assassin word is revealed or the mock is faster.
The framework also requires one metric called Quality Score corresponding to how well the task has been solved. The Quality Score, essentially a win rate, is the average number of games won (successful). The main ranking score for evaluated LLMs is the clemscore, which is the macro-average quality score multiplied by the macro-average proportion of played games to find a balance between solving most tasks and following instructions. We have also implemented the following metrics to analyse the strategies taken by models:
$\bullet$ Sensitivity: The number of revealed divided by the total team words.
$\bullet$ Efficiency: We set the bar at two target words per turn as the highest efficiency a model can reach, as that is a reasonable efficiency for humans. It is calculated as:
$min(1,\frac{1}{2}\cdot\frac{\mbox{team words revealed}}{\mbox{number of turns}})$
4.3 Evaluated Models
We evaluated open-weight and commercial models with a zero-shot setting where temp=0. We included the most recent commercial models such as: o3-mini (Jan ’25), GPT-4o (Aug ’24) Claude-3-5 (Sonnet, Oct ’24), and Gemini-2.0-Flash (Feb ’25). We also included recent open-weight models: Llama-3.1 (8B, 70B, 405B) (Grattafiori et al., 2024), Llama-3.3 (70B), Qwen2 (72B) (Yang et al., 2024), Qwen2.5 (Coder-32B, 72B, Max) (Qwen et al., 2025), and Deepseek (v3, r1) (DeepSeek-AI et al., 2024, 2025). We used the APIs of the respective commercial models. For open-weight models, we ran the inference on two NVIDIA A100 GPUs. Two Deepseek models, Llama-3.1-405B and Qwen-Max, were run via the OpenRouter API.
5 Results
Model clemscore % Played Quality Score
o3-mini 49.2 100.0 49.2
Claude-3-5 46.9 93.8 50.0
GPT-4o 45.4 93.8 48.4
Deepseek-r1 45.4 85.4 53.2
Gemini-2.0 37.7 96.2 39.2
Llama-3.1-70B 36.9 90.0 41.0
Deepseek-v3 33.8 86.9 38.9
Qwen2.5-72B 30.0 72.3 41.5
Llama-3.3-70B 29.2 80.0 36.5
Llama-3.1-405B 29.2 76.2 38.4
Qwen-max 25.4 70.0 36.3
Qwen2-72B 20.8 58.5 35.5
Qwen2.5-32B 20.8 62.3 33.3
Llama-3.1-8B 14.6 52.3 27.9
Table 1: Ranking of all benchmarked LLMs.
Experiment o3-mini GPT-4o LM-3.1 LM-3.3 Claude-3.5 Deepseek-r1 Gemini-2.0
Risk Level low 70.0 75.0 50.0 50.0 75.0 87.5 55.6
high 20.0 37.5 11.1 20.0 30.0 10.0 10.0
Association easy 100.0 100.0 100.0 100.0 100.0 100.0 100.0
difficult 20.0 10.0 20.0 0.0 20.0 28.6 12.5
Opponent none 80.0 77.8 80.0 75.0 57.1 62.5 77.8
easy 50.0 33.3 14.3 28.6 40.0 80.0 11.1
difficult 0.0 0.0 0.0 0.0 0.0 22.2 0.0
Frequency low 60.0 66.7 50.0 33.3 60.0 50.0 30.0
high 20.0 30.0 50.0 20.0 44.4 25.0 50.0
Ambiguity none 80.0 60.0 22.2 55.6 80.0 55.6 40.0
ambiguous 40.0 33.3 37.5 10.0 62.5 16.7 40.0
Concreteness concrete 80.0 50.0 66.7 44.4 50.0 88.9 40.0
abstract 20.0 60.0 0.0 16.7 40.0 50.0 40.0
Table 2: Detailed results across different experiments. Only high performing LLMs were selected. The values correspond to the Quality Score for each experiment. LM-3.1 $\rightarrow$ Llama-3.1-405B, LM-3.3 $\rightarrow$ Llama-3.3-70B
5.1 Overall Analysis
The overall results are given in Table 1 where the clemscore, Played, and Quality Score are averaged across all experiments. The first observation we make is that, as expected,larger models perform better. In line with this, commercial models outperform open-weight ones by some margin (five points between o3-mini and Deepseek-r1). o3-mini is the only model that played all episodes without once making an instruction following error in the game. However, we can see that the best model achieves only 49.2% success rate in winning the game against the mock opponent. To investigate specific experiments, we selected seven high-performing models to compare them in detail. The results are given in Table 2.
Risk level: We expected the high risk to be more complex than the low one because there are five assassin words. This expected behaviour holds for all models, e.g. o3-mini has a margin of 50 points between both experiments. In the high-risk experiment, GPT-4o achieves the best score of 37.5, which is a substantial margin of 17.5 points compared to the second-best result.
Word association: All models achieved a perfect score for the easy set. The difficult case is much more challenging as no model reaches 30 points.
Opponent level: We tested three levels of the mock opponent where the difference lies in how fast the words are revealed. The performance on the first level is significantly higher for all models as it is easier to beat the mock opponent who does not reveal any words. Even in this setting, the best models (o3-mini and Llama-3.1-405B) can only reach 80 points. However, once we switch to other levels, we see a clear drop in performance for most models, except Deepseek-r1. The difficult level shows even striking results where only Deepseek-r1 managed to achieve some performance while other models lost all episodes to the mock opponent.
Word frequency: The expectation here is that higher-frequency words are easier to play with (at least for human players). This assumption does not apply as most models are better at low frequency set, except Gemini-2.0.
Ambiguity: The expectation here is that monosemic words are easier to play with, and we can confirm that this holds for most models. Claude-3.5 is the only model to surpass 50% success rate in the ambiguous set.
Concreteness: Generally, all models perform better on the concrete set, except for GPT-4o. Interestingly, Gemini-2.0 gets equal points on both sets. It indicates that abstract words are indeed more challenging (as for humans) for models.
5.2 In-depth Analysis
Figure 2: Average number of words targeted (top) and words guessed (bottom) by models at each turn
Number of Targets, Guesses & Revealed: in Figure 2, we present the average number of words targeted and words guessed by selected models. We can see that high-performing models such as o3-mini and Deepseek-r1 generate at least 1-2 more words as targets and guesses in the beginning. Targeting and guessing more words in a single turn is the standard strategy in Codenames to win Spendlove and Ventura (2022), especially needed when playing against the mock opponent, which reveals one word at each turn. Models tend to guess fewer words than were targeted. For instance, o3-mini on average targets more than four words but guessed considerably fewer for the first turn, unlike Deepseek-r1, which targets and guesses an almost equal number of words. In Figure 7, we included the average number of target, guessed and revealed (where the guess is team word) words per model. We can see that only Deepseek-r1 exceeds the threshold of more than two words (2.2), while the rest have close values (1.5-1.9). It indicates that all models guess wrong words by revealing words from the opponent team or distractors, or even assassin words.
Success, Lose & Aborted Lose Rates:
Figure 3: Distribution of Success, Lose, Aborted episodes (up), and distribution of cases where models lose (bottom).
Figure 3 includes the distribution of episodes across Success, Lose, and Aborted. To recall, Success is when a model follows the game’s rules and beats the mock opponent by revealing the team words faster, Lose is when the mock opponent is faster or when a model reveals assassin words. Aborted is when a model does not follow formatting instructions. The top graphic shows that even best-performing models barely reach the 50% Success rates where most episodes are lost or aborted. The ratio of Aborted episodes is higher for open-weight models. The bottom graphic divides the Lost cases further into two groups: assassin word is revealed or mock opponent is faster. For most models, the main issue is losing due to being slower in revealing words than the mock opponent. Only Deepseek-r1 lost more due to revealing more assassin words than others. It shows that all models struggled with the task and lost against a strategy of revealing one word every turn.
Efficiency & Sensitivity:
Figure 4: Average efficiency and sensitivity scores for selected models.
Next, we analyse how efficient the models are regarding targeting multiple words at each turn (see metrics defined in Section 4.2). Figure 4 shows the efficiency and sensitivity scores for the selected models. We can observe that o3-mini, Deepseek-r1, Llama-3.3-70B and Deepseek-v3 have higher efficiency scores, which indicates that these models target two or more words each turn. A similar observation has also been made in Figure 2. By looking at the sensitivity scores, we can conclude that Deepseek-r1 is better at this task than Deepseek-v3 because it revealed more words (sensitivity score). Models such as Claude-3.5 and GPT-4o are more consistent (efficiency and sensitivity are closer to each other) in terms of the number of targets, guessed, and revealed words.
Model Target Halluc. Guess Halluc. Wrong #  of Guesses Guess is Clue
o3-mini 0 0 0 0
DS-r1 0 0 1 0
GPT-4o 2 3 0 0
GM-2.0 1 0 4 0
Cl-3.5 3 5 0 0
LM1-70 2 2 1 7
DS-v3 6 6 1 2
LM3-70 2 2 3 13
LM-405 10 2 16 0
QW-72 5 6 0 21
QW-M 12 8 0 15
QW-32 10 7 0 19
QW-72B 9 12 0 30
LM1-8B 3 7 18 28
Table 3: Error types and their counts for each model where an episode was aborted by the GameMaster.
Typical Errors: To understand where models fail and how higher-performing models differ from lower ones, we analysed the most common errors, then categorised them and counted each occurrence, see Table 3.
Figure 5: Transcript of an episode in Word Association “easy” experiment.
Figure 6: Transcript of an episode in Concreteness “abstract” experiment. Note that only the clue is given to player B; the list of targeted words is only to get an insight into the reasoning of player A.
The differentiating factor in high-performing models is that hallucination and instruction following issues appear more rarely. For instance, the first error type, Target Hallucinated, refers to cases where Spymaster generates a clue and targets some words, but some of these do not exist on the board (as should be known to the model). In such cases, the GameMaster aborts that episode. Similarly, Guess Hallucinated is an error that occurs on the Field Operative side where it guesses a word that does not exist on the board. Mostly, Llama-3.1-405B and Llama-3.1-8B have another issue with guessing the correct number of words that the Spymaster indicates. They tend to guess more than the number of target words (note here: models can guess less but not more than target words). Lastly, the common issue, Guess is Clue, with low-performing models is that the guessed word is the same as the clue in many cases. It shows a lack of pragmatic reasoning for choosing unrevealed candidate words from available ones on a board. In all of these cases and some minor ones, e.g. tags such as “CLUE:”, “TARGETS:”, “GUESS:” are omitted, the GameMaster aborts the game because the rules are not followed. Such instruction-following issues happen mostly with Deepseek-r1 (see Table 5).
5.3 Qualitative Analysis
We included sample outputs for the Word Association - easy experiment in Figure 6. Recall that all models achieved the perfect score for this experiment (see Table 2). The words were selected from these categories “fish”, “unit of distance”, “four footed animal”, “part of a building”, “fruit”, “an article of furniture”, “country”, “musical instrument”, “type of fuel”, “weapon”, “crime”, “sport”.
o3-mini generates clues close to the ground truth categories of words. In the second turn, it makes a slight mistake by guessing the distractor word “couch”. Similarly, GPT-4o generates similar clues but focuses on only two words at a time. An interesting case occurs with Qwen2.5-Coder-32B where, in the first turn, it targets four words with the clue “wood”. The other two models targeted the word “carp” by choosing the “fish” or “animal” categories, but Qwen2.5-Coder-32B chose the sense of “carpenter, lumber quality” to connect the clue “wood” to “carp”.
Figure 6 shows sample outputs for the Concreteness - abstract experiment. As we can see, the chosen words are not typical daily life words that would challenge human players in Codenames. Claude-3.5 manages to play this episode and win the game. We can see that it generates decent clues that combine the target words. It made one mistake by guessing a distractor word in the second turn. The gameplay by o3-mini is even more fascinating. The average number of target words is three, and it generates matching clues. However, due to the strategy of targeting and guessing more words, it gives a massive advantage to the opponent by revealing 50% of their teams’ words (“misconception”, “humbug”, “extent”, “endlessness”). Even though the model manages to reveal seven out of nine words (“heyday” and “imposition” were never revealed), it lost the game because the mock opponent revealed words (primarily due to four additional words revealed mistakenly by o3-mini).
Figure 8 includes sample episodes for the Risk level - high experiment with five assassin words. o3-mini, Claude-3.5, Gemini-2.0, Deepseek-r1 guessed one of the assassin words and lost the game. Llama3.3-70B lost the game due to guessing (six words) more than what was targeted (five words).
Figure 9 shows samples for the ambiguous words. Deepseek-v3 revealed three opponent words but still managed to win the game.
6 Discussion
Commercial vs. open: We can notice that commercial models outperform open-weight ones by some margin. We categorised the errors by models and counted them (see Table 3). The main reasons for open-weight models having a lower ratio of Played episodes are i) these models often hallucinate while choosing target words, which means they add a word in the target list that does not exist on the board, ii) hallucination also occurs by guessing words that do not exist on the board, iii) guessing the clue word itself. For instance, the performance difference between Llama3.1-70B and 405B can be explained with the bigger model: i) hallucinating target words and ii) guessing too many words.
Choice of words: The selection of words (ambiguous, abstract, high or low frequency, more assassin words) impacts the performance as expected. Of all the experiments, playing against a mock opponent that revealed two words and word associations with difficulty levels proved to be the most challenging. Similarly, abstract words seemed to be more demanding than concrete words. However, we observed that the frequency of words does not directly impact performance when looking at all model results, whereas, for humans, less frequent words might be more challenging. Similar remarks can be made for ambiguity and abstract word sets where the results are somewhat mixed and where humans are expected to find them demanding.
Reasoning models: By looking at the best performing models, we can conclude that the best of one of the commercial and open-weight options are reasoning models where Deepseek-r1 outperforming some commercial models such as Gemini-2.0 or Qwen-max. However, such an impressive performance comes at the cost of high latency. It took almost two minutes per query for r1 and two seconds for v3 (see Table 4).
Do LLMs have the required abilities to play Codenames? The models cannot play efficiently in some experiments by looking at the win rates (Quality Score) for all models. Codenames is a challenging task that involves deep language understanding, theory of mind, cooperation, and pragmatic reasoning. Our experimental results suggest that LLMs do possess knowledge about word associations, and it was shown that they can access it strategically (see Figure 6 where o3-mini targets four words with clue “animal”). Another strategy that we observed is the risk taking strategy where models target more than two words per turn to win the game (see Figure 7). Such a strategy would be a clear winner against a mock opponent that reveals only one word per turn. However, we have seen cases where this strategy resulted in actually losing the game by revealing the opponent teams’ words (see o3-mini in Figure 6). Another risky strategy was observed with the high-risk set, where models could not navigate the experiment with five assassin words. Some models still went on to target a lot of words while risking the error on the guesser side (see o3-mini on Figure 9 where it targets nine words at once and loses the game).
The experiments also reveal certain aspects of pragmatic reasoning in multi-turn tasks where if a particular clue was not utilised to guess certain target words, it has been revised (see Figure 6 where o3-mini targets the word “existentialism” with the clue “doubt” and it was not guessed, then reintroduced another clue “spirit” to the guess the same word again). The cooperation aspect can be seen where some models are consistent in terms of choosing the number of target words and how many of them were correctly guessed (see Figure 4, GPT-4o, Claude-3.5).
7 Conclusion
We implemented Codenames to benchmark LLMs by targeting their pragmatic reasoning, language understanding specifically for ad-hoc concept generation, and cooperation capabilities. We tested the most recent commercial and open-weight models on various experiments and difficulty levels. We can generally confirm that commercial models are ahead in performance compared to open-weight ones. The main reasons for better performance can be attributed to having less errors with regards to hallucinations, instruction following, and pragmatic reasoning. However, when looking at played episodes, we can say that even the best performing models do not win over 50% of the games. It clearly indicates that the task is far from being solved. Overall, the presented solution provides a clear method for benchmarking LLMs using game-based evaluation to target specific capabilities.
Limitations
The current study is restricted to only English in its current state. While we have yet to do this, translating the prompts and finding the matching word lists should be possible for other languages, too. We plan to do this in future work.
As discussed in the analysis above, some of the findings are limited to general strategies applied internally by the models. We plan to study the reasoning capabilities in detail to understand the underlying blocks that leads to certain clues or guesses to be generated.
Ethics Statement
Using paid proprietary APIs with underlying models about which little is known (training data, model architecture) in academic research is less than ideal. At the moment, the models benchmarked here seem to be the high-performing ones that are commercially used. It is our hope that more open models with high performance will be released soon, and proper research can be done with them.